OpenStudy (anonymous):

Circle O had a radius of 7.39. if m

OpenStudy (amorfide):

a diagram would help ...  AB = A²=b²+c²-2bcCos(a) A²=7.39²+7.39²-2(7.39 x 7.39) x Cos(112) A²=150.14 A=12.3 (3.S.F)